# CASteer: Cross-Attention Steering for Controllable Concept Erasure

### v3

@@ 4.1 Results
Abstract concept erasure. In this section, we present results on inappropriate content erasure based on I2P dataset DBLP:conf/cvpr/SchramowskiBDK23. I2P is a dataset of 4,703 curated prompts designed to test generative models, where most prompts lead to images containing inappropriate content. Following prior work, we test CASteer on two I2P-based tasks: 1) removing nudity, 2) removing all inappropriate content at once. For nudity removal, we utilize CASteer with steering vectors generated for the concept of "nudity". For inappropriate content removal, we use CASteer with steering vectors obtained as average of steering vectors generated for each type of inappropriate content, i.e., hate, harassment, violence, self-harm, shocking, sexual, and illegal content.
We compare our method with state-of-the-art approaches Ablating (CA) DBLP:conf/iccv/KumariZWS0Z23, FMN zhang2023forgetmenot, SLD DBLP:conf/cvpr/SchramowskiBDK23, ESD DBLP:conf/iccv/GandikotaMFB23, UCE DBLP:conf/wacv/GandikotaOBMB24, SA DBLP:conf/nips/HengS23, Receler huang2023receler, MACE DBLP:conf/cvpr/LuWLLK24, RECE DBLP:conf/eccv/GongCWCJ24, CPE lee2024cpe, AdvUn zhang2024defensive and SAeUron cywinski2025saeuron. Following prior art, we utilize the NudeNet  [^1] https://github.com/notAI-tech/NudeNet to detect nude body parts on generated images for nudity erasure task, and the NudeNet with the Q16 detector to detect inappropriate content.
@@
To assess general generation quality of CASteer, we follow prior work and run CASteer with “nudity” steering vectors on prompts from COCO-30k DBLP:conf/eccv/LinMBHPRDZ14. We report FID DBLP:conf/nips/HeuselRUNH17 for general visual quality and CLIP score DBLP:conf/emnlp/HesselHFBC21 for image-prompt alignment. Results are presented in Tab. 3. Both versions of CASteer have better FID than prior art.
@@
Method
Locality
CLIP-30K($\uparrow$)
FID-30K($\downarrow$)
@@
Figure 2: SPM failure in removing implicitly defined concepts (SD-1.4). Top: CASteer, Bottom: SPM. Left: “a mouse from Disneyland,” Right: “a yellow Pokemon.” CASteer erases Mickey and Pikachu concepts despite not being explicitly named, while SPM fails.
@@
Concrete concepts erasure. To assess ability of CASteer to remove concrete concepts, we follow the experimental setup of SPM DBLP:conf/cvpr/Lyu0HCJ00HD24. In this setting, the concept to be erased is Snoopy, and images of five additional concepts (Mickey, Spongebob, Pikachu, dog and legislator) are generated to test the capability of the method to preserve content not related to the concept being removed. The first four of these are specifically chosen to be semantically close to the concept being removed to show the model’s ability to perform precise erasure. Following SPM DBLP:conf/cvpr/Lyu0HCJ00HD24, we augment each concept using $80$ CLIP DBLP:conf/icml/RadfordKHRGASAM21 templates, and generate $10$ for each concept-template pair, so that for each concept there are $800$ images. We evaluate the results using two metrics. First, we utilize CLIP Score (CS) DBLP:conf/emnlp/HesselHFBC21 to confirm the level of the existence of the concept within the generated content. Second, we calculate FID DBLP:conf/nips/HeuselRUNH17 scores between the set of original generations of SD-1.4 model and a set of generations of the steered model. We use it to assess how much images of additional (non-Snoopy) concepts generated by the steered model differ from those of generated by the original model. A higher FID value demonstrates more severe generation alteration. We present the results in Tab. 15. In Fig. 3, we also show two types of plots. Fig. 3(a) pictures normalized clip score of source concept, i.e. “Snoopy” (the lower the better) versus mean normalized clip scores of other concepts (the higher the better). Normalization is done to ensure equal importance of all the concepts in the mean, and done by dividing clip score of images produced by erasing method by clip score of images produced by vanilla SD-1.4. Methods on the left of the plot erase Snoopy well, and methods on top of the plot preserve other concepts well. Fig. 3(b) pictures normalized clip score of source concept versus mean FID scores of other concepts (the lower the better). Methods on the left of the plot erase Snoopy well, and methods on top of the plot tend not to affect images of other concepts much.
@@
Results show that CASteer maintains good balance between erasing unwanted concept, while preserving other concepts intact. ESD  DBLP:conf/iccv/GandikotaMFB23 and Receler erase Snoopy well, but also highly affect other concepts, especially related ones such as Mickey or Spongebob. Note that their CS of unrelated concepts (e.g. “Mickey” or “Spongebob”) are significantly lower than that of original SD-1.4, indicating that these concepts are also being affected when erasure of “Snoopy” is done. High FID of these methods on these concepts supports this observation. SAFREE shows a reduced level of Snoopy erasure compared to that of CASteer, and has lower CS and higher FID on all the other concepts. SPM keeps unrelated concepts almost intact (High CS and low FID), but has a much lower intensity of Snoopy erasure. Moreover, SPM fails to erase implicitly defined concepts (see Fig. 2 and Sec. D in supplementary). We provide qualitative results on comparisons in the supplementary.
@@
CASteer is capable of erasing implicitly defined concepts. We check what happens if we define the prompts implicitly, e.g., “A mouse from Disneyland”. We run CASteer and SPM trained on the Mickey concept on these prompts and show the results in Fig. 2. We clearly see that SPM fails to erase the concepts when they are not explicitly defined. In contrast, our method does a much better job of erasing the concepts, despite being implicitly defined. This is also supported by the results on nudity erasure in Tab. 1, as considered datasets contain specially selected adversarial nudity prompts. We provide additional results on implicitly defined prompts in the supplementary.
Overall experimental results show that CASteer performs precise erasure of both concrete and abstract concepts and concepts defined implicitly while leaving other concepts intact and not affecting the overall quality of generated images. More qualitative results showing the performance of CASteer on prompts related and not related to the target concept can be found in the supplementary.
@@ 4.2 Ablation Study
Steering other layers. As mentioned in Sec. 3.1, we ablate to determine for which type of layer in the DiT backbone steering is most effective. We show in the supplementary that steering the CA outputs is the most effective. We also ablate steering only a fraction of CA layers in Sec. G.4.
Number of prompt pairs to construct steering vectors. In the supplementary, we provide an ablation on the number of prompt pairs needed to produce high-quality outputs after steering. We find that as little as 50 prompts is enough for the steering vectors to capture the desired concept well.
Interpretation of steering vectors. Here, we propose a way to interpret the meaning of steering vectors generated by CASteer. Suppose we have steering vectors generated for a concept $X$ $\{ca^{X}_{it}\},1\leqslant i\leqslant l,1\leqslant t\leqslant T$, where $l$ is the number of model layers and $T$ is the number of denoising steps performed for generating steering vectors. To interpret these vectors, we prompt the diffusion model with a placeholder prompt “X" and at each denoising step, we substitute outputs of the model’s CA layers with corresponding steering vectors. This conditions the diffusion model only on the information from the steering vectors, completely suppressing other information from the text prompt. Results are presented in Fig. 38 and in the supplementary.
UMap. We generate steering vectors for all vocabulary tokens of SDXL text encoders and apply UMap DBLP:journals/corr/abs-1802-03426 on these steering vectors. We present the results in the appendix, showing that structure emerges in the space of these steering vectors, similar to that of Word2Vec DBLP:journals/corr/abs-1301-3781, supporting that steering vectors carry the meaning of the desired concept.
Modern models. We show qualitative results in SANA and SDXL in Fig. 4. We provide more qualitative and quantitative results in SDXL (Sec. E) and SANA (Sec. F).
@@ C.2 Prompts for human-related concepts
Following Receler, we use $\mathbf{e}=$“nudity" for nudity erasure, and set of $\{``hate",``harassment",``violence",``suffering",``humiliation",``harm",``suicide",``sexual",``nudity",``bodilyfluids",``blood$"\} for harmful content erasure.
@@ C.3 Prompts for style manipulation
These prompts are used in Sec. H.3
For erasing or adding style, we use prompt pairs of the form:
@@ Appendix D User studies
Lastly, we want to see how well SPM, Receeler and CASteer preserve concepts“Mickey" when “Snoopy" is removed. We provide users with 3 images, one for each model, and ask them to rank these images based on how much the “Mickey" concept is preserved in the images. We thus ask them “Rank the images in order from higher level of “Mickey" concept present (1) to lowest level of "Mickey" concept present (3).". We give each user a randomly selected subset of 15 triplets of images. We can observe in Tab. 7 that SPM preserves the best the concept of Mickey, followed by CASteer. However, from Tab. 6 we see that SPM has a much higher tendency to preserve concepts, even the ones that should be removed, such as “Snoopy". Our model has been considered by users a more reliable model for both removing some concepts and preserving the rest.
@@
Appendix E Results on SDXL model
@@ E.1 Experimental setup
E.2 Quantitative results
We use the same experimental setups as for SD-1.4, described in Sec. 4. Results are shown in Tab.8910,11.
Table 8: Quantitative results on nudity removal based on I2P DBLP:conf/cvpr/SchramowskiBDK23 dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@ E.2 Quantitative results
Table 9: Quantitative results on inappropriate content removal based on I2PDBLP:conf/cvpr/SchramowskiBDK23 dataset. Detection of inappropriate content is done by Q16 schramowski2022can.
@@
Method
Locality
CLIP-30K($\uparrow$)
FID-30K($\downarrow$)
@@
E.3 Qualitative results
@@
We see that our method removes Snoopy well (see fig. 5 while preserving other concepts well (see fig.  6, 7, 8, 9, 10). In fact, most of the images of non-related concepts generated with CASteer applied are almost identical to those generated by vanilla SDXL.
Second, we show images generated on COCO-30k prompts with applied CASteer for nudity removal. We see that quality of generated images does not degrade, supporting quantitative results of Tab.10
Figure 5: Images generated with “Snoopy" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 6: Images generated with “Mickey" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 7: Images generated with “Spongebob" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 8: Images generated with “Pikachu" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 9: Images generated with “dog" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 10: Images generated with “legislator" prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy".
Figure 11: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Figure 12: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Figure 13: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Figure 14: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity".
Appendix F Results on SANA model
F.1 Experimental setup

### v4

@@ 4.1 Results
Abstract concept erasure. In this section, we present results on inappropriate content erasure based on I2P dataset Schramowski et al. (2023). I2P is a dataset of 4,703 curated prompts designed to test generative models, where most prompts lead to images containing inappropriate content. Following prior work, we test CASteer on two I2P-based tasks: 1) removing nudity, 2) removing all inappropriate content at once. For nudity removal, we utilize CASteer with steering vectors generated for the concept of "nudity". For inappropriate content removal, we use CASteer with steering vectors obtained as average of steering vectors generated for each type of inappropriate content, i.e., hate, harassment, violence, self-harm, shocking, sexual, and illegal content.
We compare our method with the following state-of-the-art approaches: Ablating (CA) Kumari et al. (2023), FMN Zhang et al. (2024a), SAeUron Cywiński and Deja (2025) SLD Schramowski et al. (2023), ESD Gandikota et al. (2023), UCE Gandikota et al. (2024), MACE Lu et al. (2024), SA Heng and Soh (2023), Receler Huang et al. (2024), RECE Gong et al. (2024), CPE Lee et al. (2025), AdvUn Zhang et al. (2024c) and . Following prior art, we utilize the NudeNet  [^1] https://github.com/notAI-tech/NudeNet to detect nude body parts on generated images for nudity erasure task, and the NudeNet with the Q16 detector to detect inappropriate content.
@@
To assess general generation quality of CASteer, we follow prior work and run CASteer with “nudity” steering vectors on prompts from COCO-30k Lin et al. (2014). We report FID Heusel et al. (2017) for general visual quality and CLIP score Hessel et al. (2021) for image-prompt alignment. Results are presented in Tab. 3. Both versions of CASteer have better FID than prior art.
@@
Method Locality
CLIP-30K($\uparrow$) FID-30K($\downarrow$)
@@
Figure 3: SPM and DoCo failure in removing implicitly defined concepts (SD-1.4). Top: CASteer, Middle: SPM, Bottom: DoCo. Prompt: “a mouse from Disneyland,” CASteer erases Mickey concepts despite not being explicitly named, while SPM and DoCo fail.
@@
Concrete concepts erasure. To assess ability of CASteer to remove concrete concepts, we follow the experimental setup of SPM Lyu et al. (2024). In this setting, the concept to be erased is Snoopy, and images of five additional concepts (Mickey, Spongebob, Pikachu, dog and legislator) are generated to test the capability of the method to preserve content not related to the concept being removed. The first four of these are specifically chosen to be semantically close to the concept being removed to show the model’s ability to perform precise erasure. Following SPM Lyu et al. (2024) and DoCo Wu et al. (2025), we augment each concept using $80$ CLIP Radford et al. (2021) templates, and generate $10$ for each concept-template pair, so that for each concept there are $800$ images. We evaluate the results using two metrics. First, we utilize CLIP Score (CS) Hessel et al. (2021) to confirm the level of the existence of the concept within the generated content. Second, we calculate FID Heusel et al. (2017) scores between the set of original generations of SD-1.4 model and a set of generations of the steered model. We use it to assess how much images of additional (non-Snoopy) concepts generated by the steered model differ from those of generated by the original model. A higher FID value demonstrates more severe generation alteration. We present the results in Tab. 17. In Fig. 4, we also show two types of plots. Fig. 4(a) pictures normalized clip score of source concept, i.e. “Snoopy” (the lower the better) versus mean normalized clip scores of other concepts (the higher the better). Normalization is done to ensure equal importance of all the concepts in the mean, and done by dividing clip score of images produced by erasing method by clip score of images produced by vanilla SD-1.4. Methods on the left of the plot erase Snoopy well, and methods on top of the plot preserve other concepts well. Fig. 4(b) pictures normalized clip score of source concept versus mean FID scores of other concepts (the lower the better). Methods on the left of the plot erase Snoopy well, and methods on top of the plot tend not to affect images of other concepts much.
@@
Results show that CASteer maintains good balance between erasing unwanted concept, while preserving other concepts intact. ESD  Gandikota et al. (2023) and Receler erase Snoopy well, but also highly affect other concepts, especially related ones such as Mickey or Spongebob. Note that their CS of unrelated concepts (e.g. “Mickey” or “Spongebob”) are significantly lower than that of original SD-1.4, indicating that these concepts are also being affected when erasure of “Snoopy” is done. High FID of these methods on these concepts supports this observation. SAFREE shows a reduced level of Snoopy erasure compared to that of CASteer, and has lower CS and higher FID on all the other concepts. SPM keeps unrelated concepts almost intact (High CS and low FID), but has a much lower intensity of Snoopy erasure. Moreover, SPM fails to erase implicitly defined concepts (see Fig. 3 and Sec. D in the appendix). We provide qualitative results on comparisons in Fig. 5 and in the appendix.
@@
Style erasure. Following SAFREE Yoon et al. (2024) and ESD Gandikota et al. (2023), we also evaluate our method on two artist-style removal tasks. One task focuses on the styles of five famous artists (Van Gogh, Picasso, Rembrandt, Warhol, Caravaggio) and the other uses five modern artists (McKernan, Kinkade, Edlin, Eng, Ajin: Demi-Human), with the task being removing the style of Van Gogh and McKernan. Following SAFREE Yoon et al. (2024), we use LPIPS Zhang et al. (2018) and prompt GPT-4o to identify an artist on generated images as evaluation metrics.
Table 4: Comparison of Artist Concept Removal tasks: Famous (left) and Modern artists (right).
Remove “Van Gogh” Remove “Kelly McKernan”
Method $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$ $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$
SD-v1.4 - - 0.95 0.95 - - 0.80 0.83
CA (Kumari et al., 2023) 0.30 0.13 0.65 0.90 0.22 0.17 0.50 0.76
RECE (Gong et al., 2024) 0.31 0.08 0.80 0.93 0.29 0.04 0.55 0.76
UCE (Gandikota et al., 2024) 0.25 0.05 0.95 0.98 0.25 0.03 0.80 0.81
SLD-Medium (Schramowski et al., 2023) 0.21 0.10 0.95 0.91 0.22 0.18 0.50 0.79
SAFREE (Yoon et al., 2024) 0.42 0.31 0.35 0.85 0.40 0.39 0.40 0.78
Ours (w/o clip) 0.46 0.31 0.35 0.88 0.54 0.27 0.05 0.80
Ours (clip) 0.44 0.30 0.25 0.86 0.54 0.28 0.05 0.81
We follow SAFREE Yoon et al. (2024) evaluation procedure, please refer to it or our appendix (Sec. E) for more details on the procedure and metrics. From Tab. 4, we see that CASteer shows the best results in style removal (see columns LPIPSe and Acce), while preserving other styles well (see columns LPIPSu and Accu). Among all the approaches, CASteer achieves great balance between target style removal and preservation of other styles.
CASteer is capable of erasing implicitly defined concepts. We check what happens if we define the prompts implicitly, e.g., “A mouse from Disneyland”. We run CASteer, SPM and DoCo trained on the Mickey concept on these prompts and show the results in Fig. 3. We clearly see that SPM and DoCo fail to erase the concepts when they are not explicitly defined. In contrast, our method does a much better job of erasing the concepts, despite being implicitly defined. This is also supported by the results on nudity erasure in Tab. 1, as considered datasets contain specially selected adversarial nudity prompts. We provide additional results on implicitly defined prompts in the appendix.
Overall experimental results show that CASteer performs precise erasure of both concrete and abstract concepts and concepts defined implicitly while leaving other concepts intact and not affecting the overall quality of generated images. More qualitative results showing the performance of CASteer on prompts related and not related to the target concept can be found in the appendix.
@@
Figure 6: Qualitative results on SANA (right) on removing style of “Van Gogh”. Top: original model generations, bottom: generations of model steered to remove “Van Gogh” style. Left 2 images correspond to prompts with “Van Gogh style”, images on the right correspond to prompts mentioning other artists ( “Picasso”, “Rembrandt”, “Warhol”, “Caravaggio”)
@@ 4.2 Ablation Study
Steering other layers. As mentioned in Sec. 3.1, we ablate to determine for which type of layer in the DiT backbone steering is most effective. We show in the appendix that steering the CA outputs is the most effective. We also ablate steering only a fraction of CA layers in Sec. H.5.
Number of prompt pairs to construct steering vectors. In the appendix, we provide an ablation on the number of prompt pairs needed to produce high-quality outputs after steering. We find that as little as 50 prompts is enough for the steering vectors to capture the desired concept well.
Interpretation of steering vectors. Here, we propose a way to interpret the meaning of steering vectors generated by CASteer. Suppose we have steering vectors generated for a concept $X$ $\{ca^{X}_{it}\},1\leqslant i\leqslant l,1\leqslant t\leqslant T$, where $l$ is the number of model layers and $T$ is the number of denoising steps performed for generating steering vectors. To interpret these vectors, we prompt the diffusion model with a placeholder prompt “X” and at each denoising step, we substitute outputs of the model’s CA layers with corresponding steering vectors. This conditions the diffusion model only on the information from the steering vectors, completely suppressing other information from the text prompt. Results are presented in Fig. 41 and in the appendix.
UMap. We generate steering vectors for all vocabulary tokens of SDXL text encoders and apply UMap McInnes and Healy (2018) on these steering vectors. We present the results in the appendix, showing that structure emerges in the space of these steering vectors, similar to that of Word2Vec Mikolov et al. (2013), supporting that steering vectors carry the meaning of the desired concept.
Modern models. We show qualitative results in SANA and SDXL in Fig. 6. We provide more qualitative and quantitative results in SDXL (Sec. F) and SANA (Sec. G).
@@ C.2 Prompts for human-related concepts
Following Receler, we use $\mathbf{e}=$“nudity” for nudity erasure, and set of $\{``hate",``harassment",``violence",``suffering",``humiliation",``harm",``suicide",\allowbreak``sexual",``nudity",``bodilyfluids",``blood$"\} for harmful content erasure.
@@ C.3 Prompts for style manipulation
For erasing or adding (Sec. I.3) style, we use prompt pairs of the form:
@@ Appendix D User studies
Lastly, we want to see how well SPM, Receeler and CASteer preserve concepts “Mickey” when “Snoopy” is removed. We provide users with 3 images, one for each model, and ask them to rank these images based on how much the “Mickey" concept is preserved in the images. We thus ask them “Rank the images in order from higher level of “Mickey” concept present (1) to lowest level of “Mickey” concept present (3).". We give each user a randomly selected subset of 15 triplets of images. We can observe in Tab. 8 that SPM preserves the best the concept of Mickey, followed by CASteer. However, from Tab. 7 we see that SPM has a much higher tendency to preserve concepts, even the ones that should be removed, such as “Snoopy”. Our model has been considered by users a more reliable model for both removing some concepts and preserving the rest.
@@
Appendix E Details on style removal task
The task focuses on the styles of five famous artists (Van Gogh, Picasso, Rembrandt, Warhol, Caravaggio) and the other uses five modern artists (McKernan, Kinkade, Edlin, Eng, Ajin: Demi Human). Following SAFREE, we use $LPIPS$ and prompt GPT-4o to identify an artist on generated images as evaluation metrics.
Following recent works, we experiment with removing one style of famous artist (Van Gogh) and one style of a modern artist (McKernan). We assess both how well our methods remove the desired style (e.g., Van Gogh) and preserve other styles in the same batch (e.g. Picasso, Rembrandt, Warhol and Caravaggio, if we remove Van Gogh). $LPIPS_{e}$ shows $LPIPS$ values for the images generated, when prompted with target style (“Van Gogh” or “Kelly McKernan”), and $LPIPS_{u}$, shows $LPIPS$ values for the images generated with the other 4 styles. $Acc_{e}$ shows the accuracy of GPT-4o answers when asked to identify an artist on images generated with target style in prompt (“Van Gogh” or “Kelly McKernan”), and $Acc_{u}$ shows the accuracy of GPT-4o answers when asked to identify an artist on images generated with other styles in prompt. The goal of any style removal method is to lower $LPIPS_{e}$ and $Acce$ (i.e., successfully remove target style) while maintaining $LPIPS_{u}$ and $Acc_{u}$ high (i.e., not affecting the generation of other styles).
Questions to GPT-4o are formulated as “Is this picture in s style? Just tell me Yes or No.”, where $s$ represents style, e.g., s =“Andy Warhol” or s =“Van Gogh”. The number of “Yes” answers divided by the number of total answers is considered as Acc metric. We report results averaged over 3 runs for this metric.
Appendix F Results on SDXL model
@@ F.1 Experimental setup
F.2 Quantitative results
We use the same experimental setups as for SD-1.4, described in Sec. 4. Results are shown in Tab. 910111213.
Table 9: Quantitative results on nudity removal based on I2P Schramowski et al. (2023) dataset. Detection of nude body parts is done by Nudenet at a threshold of 0.6. F: Female, M: Male. The best results are highlighted in bold, second-best are underlined.
@@ F.2 Quantitative results
Table 10: Quantitative results on inappropriate content removal based on I2PSchramowski et al. (2023) dataset. Detection of inappropriate content is done by Q16 Schramowski et al. (2022).
@@
Method Locality
CLIP-30K($\uparrow$) FID-30K($\downarrow$)
@@
Table 13: Comparison of Artist Concept Removal tasks on SDXL model: Famous (left) and Modern artists (right).
Remove “Van Gogh" Remove “Kelly McKernan"
Method $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$ $_{e}\uparrow LPIPS$$_{u}\downarrow$ $_{e}\downarrow Acc$$_{u}\uparrow$
SDXL - - 0.95 0.95 - - 1.00 0.86
Ours (w/o clip) 0.34 0.19 0.00 0.84 0.39 0.12 0.00 0.79
Ours (clip) 0.35 0.19 0.00 0.85 0.39 0.12 0.05 0.83
F.3 Qualitative results
@@
First, we show results on removing “Snoopy” concept when generating images with four prompt templates: “An origami XX”, “A drawing of the XX”, “A photo of a cool XX” and “An art of the XX”, where $X\in$ [“Snoopy”, “Mickey”, “Spongebob”, “Pikachu”, “dog", “legislator”]. CASteer is applied with removal strength $\beta=2$.
We see that our method removes Snoopy well (see Fig. 7 while preserving other concepts well (see Fig.  8,  9,  10,  11,  12). In fact, most of the images of non-related concepts generated with CASteer applied are almost identical to those generated by vanilla SDXL.
Second, we show images generated on COCO-30k prompts with applied CASteer for nudity removal. We see that quality of generated images does not degrade, supporting quantitative results of Tab. 11 (see Fig. 131415)
Next, on Fig. 16 and Fig. 17 we show examples of CASteer applied for style removal.
Figure 7: Images generated with “Snoopy” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 8: Images generated with “Mickey” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 9: Images generated with “Spongebob” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 10: Images generated with “Pikachu” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 11: Images generated with “dog” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 12: Images generated with “legislator” prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “Snoopy”.
Figure 13: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity”.
Figure 14: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity”.
Figure 15: Images generated with different COCO-30k prompts with different seeds. Top: original SDXL, bottom: CASteer applied for removing the concept of “nudity”.
Figure 16: Examples of style removal for SDXL. Top: original SDXL, bottom: CASteer applied for removing the style of “Van Gogh
Figure 17: Examples of style removal for SDXL. Top: original SDXL, bottom: CASteer applied for removing the style of “mosaic style” (left 4 images) and “origami style” (right 4 images)
Appendix G Results on SANA model
G.1 Experimental setup